# How and Why LLMs Use Deprecated APIs in Code Completion? An Empirical Study

### v2

@@ Abstract
To address this gap, we conducted the first evaluation study on deprecated API usage in LLM-based code completion. This study involved seven advanced LLMs, 145 API mappings from eight popular Python libraries, and 28,125 completion prompts. The study results reveal the status quo and root causes of deprecated API usage in LLM-based code completion from the perspectives of model, prompt, and library. Based on these findings, we propose two lightweight fixing approaches, ReplaceAPI and InsertPrompt, which can serve as baseline approaches for future research on mitigating deprecated API usage in LLM-based completion. Additionally, we provide implications for future research on integrating library evolution with LLM-driven software development.
@@ I Introduction
Large language models (LLMs) [1, 2, 3, 4, 5, 6] have significantly advanced various aspects of software engineering, including code completion [7, 8], code understanding [9], and program repair [10, 11]. These models, pre-trained or fine-tuned with extensive knowledge of code on large corpora, are effective for tailoring to different downstream tasks. In the realm of code completion, the state-of-the-art has evolved from statistics-based methods [12, 13] to LLM-based techniques [14, 15, 16]. Code completion is a sophisticated task that suggests variables, functions, classes, methods, and even entire code blocks, which depends both on tools’ capability and developers’ practical needs.
@@
Motivation. To accelerate development, developers heavily rely on third-party libraries, interacting with them through Application Programming Interfaces (APIs). However, this reliance presents a challenge for code completion tools. Third-party libraries constantly evolve to undergo refactorings [17], fix bugs [18], apply security patches [19], or introduce new features. This rapid evolution leads to frequent API changes, with older APIs being deprecated and replaced by newer ones. Deprecated APIs are discouraged to use because they might not work well with newer features or data. These outdated APIs will eventually disappear in future library updates [20]. Taking PyTorch [21], a popular deep learning library for instance, the API torch.gels() was deprecated in version 1.2 (August 2019) in favor of torch.lstsq(). Then, torch.lstsq() was deprecated in version 1.9 (June 2021) in favor of torch.linalg.lstsq(). Consequently, newly developed code should avoid using the deprecated torch.gels() and torch.lstsq(). Therefore, it’s crucial for code completion tools to suggest the correct, up-to-dated APIs to developers.
@@
Literature. However, to the best of our knowledge, the capabilities of LLM-based code completion regarding API deprecation is understudied [22, 23]. Although there emerges a substantial number of evaluation on code completion, a body of research focused on assessing the overall accuracy across various benchmarks [9, 24, 25, 26, 27]. Interestingly, Ding et al. [28] identified undefined names and unused variables as the most common syntactic errors produced by LLMs in Python code completions. Izadi et al. [29] found that incorrect function name predictions were prevalent, accounting for 23% of all token-level errors. Furthermore, Liu et al. [30] highlighted the issue of hallucinations in LLM-generated code. Their findings indicate the prevalence and potential risks of using unexpected APIs. Nevertheless, while researchers have noted the prevalence of incorrect function name predictions, they have not investigated this issue in depth. Library APIs, which constitute an important part in predicting external function names, are worth attached importance to.
@@
Study. To address this gap, we conducted a study to examine the issue of deprecated API usage in LLM-based code completion. The study aims to answer the primary research question:
@@
What are the status quo and root causes of deprecated and replacing API usage in LLM-based code completion?
@@
Model Perspective (RQ1) investigates the status quo and root causes based on the performance of various LLMs;
@@
Prompt Perspective (RQ2) examines the impact of different prompts on the status quo and root causes;
@@
Library Perspective (RQ3) analyzes the status quo and root causes across different libraries.
@@
To address these research questions, we conducted a series of experiments involving various libraries and LLMs. We collected 145 API mappings between deprecated APIs and their replacements from eight popular Python libraries. Based on these mappings, we retrieved 9,022 outdated functions and 19,103 up-to-dated functions using the deprecated APIs and replacing APIs, respectively. Each outdated or up-to-dated function was transformed into a line-level completion prompt by identifying the deprecated or replacing API and removing the containing and subsequent lines. The located API is referred to as the reference API. These prompts were then inputted into seven advanced code LLMs, including CodeLlama [5] and GPT-3.5, to generate completions and analyze the predicted API usages. If the predicted API usage corresponds to either the deprecated or replacement version of the reference API, it is annotated as plausible; otherwise, it is annotated as irrelevant.
@@
The study results reveal the following findings: Answer to RQ1: All seven evaluated LLMs encounter challenges in predicting plausible API usages and face issues with deprecated API usages, due to the presence of deprecated API usages during model training and the absence of API deprecation knowledge during model inference. Answer to RQ2: For the two categories of prompts derived from outdated and up-to-dated functions, the LLMs’ performance in predicting plausible and deprecated API usages differs significantly, influenced by the distinct code context characteristics of these prompts. Answer to RQ3: Across the eight libraries, the LLMs exhibit significant differences in their use of deprecated APIs, influenced by the characteristics of API deprecations during the evolution of these libraries.
@@
Approach. Based on the study results and findings, we developed two lightweight fixing approaches to mitigate deprecated API usage in LLM-based code completion. Given a completion containing a deprecated API usage, the first approach, named ReplaceAPI, directly replaces the deprecated API usage with the replacement and regenerates the remaining parts (e.g., argument list) during the decoding process. The second approach, named InsertPrompt, inserts an additional replacing prompt after the original prompt to guide the LLMs to use the replacement API and then regenerate the completions. We then evaluate the effectiveness of the proposed approaches in terms of fixing deprecated API usages and the accuracy in predicting line-level completions (RQ4). The evaluation results demonstrate that ReplaceAPI effectively addresses deprecated API usages for all evaluated open-source LLMs, achieving fix rates exceeding 85% with acceptable accuracy measured by Edit Similarity and Exact Match compared to ground-truth completions. While InsertPrompt does not currently achieve sufficient effectiveness and accuracy in fixing completions containing deprecated API usages, it shows potential for future exploration.
To summarize, this paper makes the following contributions:
  • The first study that reveals the status quo and root causes of deprecated API usages from model perspective, prompt perspective, and library perspective. The study involves seven advanced LLMs, 145 API mappings from eight popular Python libraries, and 28,125 prompts derived from 9,022 outdated functions and 19,103 up-to-dated functions.
@@
  • Two lightweight fixing approaches, named ReplaceAPI and InsertPrompt, which can serve as baseline approaches for future research on mitigating deprecated API usage in LLM-based completion.
  • The implications that provide potential research directions on the synergy of library evolution and LLM-driven software development.
@@ II Related Work
We review the related work with respect to library evolution and LLM-based code completion.
@@ II-A Library Evolution
Library evolution involves refactorings [17], bug fixes [18], and new feature introductions. Typically, refactorings can deprecate old APIs and introduce new replacements. Several studies have examined the reasons that developers deprecate APIs and how the clients react to such deprecations [31, 32, 33, 34]. The reasons include improving readability, reducing redundancy, avoiding bad code practices and fixing functional bugs. Deprecated APIs can affect hundreds of clients [35], particularly when clients struggle to keep pace with rapidly evolving software [36]. McDonnell et al. [37] found that only 22% of outdated API usages are eventually upgraded to use replacement APIs. Similarly, Hora et al. [38] found that client developers consumes considerate time to discover and apply replacing APIs, with the majority of systems not reacting at all. When clients do not upgrade their APIs, they silently accumulate technical debt in the form of future API changes when they finally upgrade [39]. To locate the replacing API, existing works leverage change rules written by developers [40], developer recordings [41], similarity matching [42], mining API usage in libraries [43], and in client projects [44]. Henkel and Diwan [41] developed an IDE plugin that allows library developers to record API refactoring actions and client developers to replay them. Godfrey and Zou [45] proposed a semi-automated origin analysis using similarities in name, declaration, complexity metrics, and call dependencies. Wu et al. [46] introduced a hybrid approach combining call dependency and text similarity analysis to identify API change rules. Recently, Huang2021 [47] proposed RepFinder to find replacement APIs for deprecated APIs in library updates from multiple sources.
In this work, we aim to comprehend the statuses and causes of deprecated API usages in LLM-based code completion and provide implications for mitigating the deprecated API usages.
@@ II-B LLM-based Code Completion
Code completion is an important functionality in modern IDEs and editors. Historically, researchers have explored statistical models [12, 13]. With the advent in natural language processing, researchers have embraced deep learning for code completion [7, 8] because they are similar in token-based prediction. To explore the capability of code completion tools driven by large language models (LLMs) [14, 15, 16], numerous evaluations of LLMs have been proposed. Ciniselli et al. [26, 27] conducted a large-scale study exploring the accuracies of state-of-the-art Transformer-based models in supporting code completion at various granularity levels, from single tokens to entire code segments. Zeng et al. [9] found that pre-trained models significantly outperform non-pre-trained state-of-the-art techniques in program understanding tasks. They also reveal that no single pre-trained model dominates across all tasks. Xu et al. [24] evaluated the performance of LLMs on the HumanEval dataset. Ding et al. [28] identified undefined names and unused variables as the most common errors produced by language models in Python code completions. Izadi et al. [29] evaluated the LLMs using real auto-completion usage data across 12 languages. They found that incorrect function name predictions, were prevalent, accounting for 23% of all token-level errors errors. Besides, Liu et al. [25] proposed EvaluPlus, which benchmarks the functional correctness of LLM-synthesized code using test cases. In addition to accuracy concerns, LLM-based approaches face issues such as security vulnerabilities and hallucinations. Sallou et al. [48] explored threats posed by LLMs, including unpredictability in model evolution, data leakage, and reproducibility. Liu et al. [30] categorized the hallucinations brought by LLM-generated code. More notably, Wu et al. [49] introduced VersiCode, the pioneering dataset aimed at evaluating the capability of large language models in generating verifiable code tailored to specific library versions.
The findings on incorrect function predictions partially motivate our study. However, our study focuses on the severity of predicting deprecated API usages in LLM-based code completion.
@@ III Study Setup
We chose Python, a popular programming language which ranks first among the most popular ones based in the recent year [50]. We targeted eight popular Python libraries. Five of these libraries were used in a previous study on Python API deprecation [19], including Numpy, Pandas, scikit-learn, SciPy, and seaborn. Additionally, we added three popular deep learning libraries, i.e., TensorFlow [^1] https://www.tensorflow.org/, PyTorch [^2] https://pytorch.org/, and Transformers [^3] https://huggingface.co/docs/transformers/index. The setup of our study is presented in Figure 1. It includes four steps. The API Mapping Collection step gathers mappings between deprecated APIs and their replacements from various libraries. The Completion Prompt Construction step involves creating completion prompts by identifying instances of deprecated and replacement API usage in open-source Python repositories. The LLM-Based Code Completion step uses various LLMs to generate code completions for these prompts. Finally, the Completion Result Annotation step automatically annotates the generated completions and calculates relevant metrics.
@@ III-A API Mapping Collection
We identified API mappings (i.e., deprecated APIs and the mapping replacements) from the documentation and change logs from each library following the previous study [19]. Specifically, we reviewed the documentation and change logs of each library and manually look for deprecated API occurrences which indicate the corresponding the mapping replacements. For instance, in the API documentation of PyTorch, version 1.9.0 [51], a deprecation message indicates that “torch.lstsq() is deprecated in favor of torch.linalg.lstsq() and will be removed in a future PyTorch release.”, where the mapping of the deprecated API to the replacing API is torch.lstsq $\to$ torch.linalg.lstsq. For one-to-many mappings (i.e., one deprecated API mapped to multiple replacing APIs), we split them into many one-to-one mappings. In total, the authors obtained 145 API mappings. The statistics of the API mappings are presented Table I.
TABLE I: Statistics of our Collected API Mappings in Eight Python Libraries
  Library Version # Mappings # Functions
@@
  Numpy 1.26.4 3 567 2,988
Pandas 2.2.2 10 69 69
scikit-learn 1.5.0 18 985 1,197
SciPy 1.13.0 4 245 1,458
seaborn 0.13.2 3 904 1,329
TensorFlow 2.16.1 57 1,491 4,830
PyTorch 2.3.0 21 4,726 6,406
Transformers 4.40.2 29 100 63
@@ III-B Completion Prompt Construction
We constructed code completion prompts by searching the deprecated API and replacing API usages from open-source Python repositories.
@@
We utilized Sourcegraph [52], a widely used code search service. It supports integration with GitHub where we can retrieve Python source files from millions of open-source code repositories. For each deprecated or replacing API, we constructed search queries using both its full qualified name (FQN) (e.g., torch.lstsq) and a logical disjunction of its constituent parts (e.g., “torch AND linalg AND lstsq”) to ensure comprehensive retrieval. For each retrieved Python source file, we parsed it into an Abstract Syntax Tree (AST) and extracted the containing functions that invoked the deprecated or replacing APIs. Specifically, we located function definition nodes in the AST and traversed its descendants. For each descendant, we checked if it is a function call node and matched the function call to the deprecated or replacing APIs. To correctly match the function call via API FQNs, we performed lightweight object type resolution and API alias resolution, similar to [19].
@@
  • API Alias Resolution: Developers can alias packages, classes, and functions in Python using the import-as feature [53]. This mechanism requires resolving API aliases by analyzing import statements. For example, the pandas package is often imported with the alias “pd” via the statement import pandas as pd. In this case, pd.DataFrame.loc() was resolved to pandas.DataFrame.loc(). Additionally, Python provides a from-import mechanism allowing developers to use APIs with short names instead of their FQNs. For example, through from torch.linalg import lstsq, the API in torch.linalg can be directly called via lstsq(). These short names were also resolved by analyzing the import statements.
After the lightweight object type resolution and API alias resolution, we obtained the corresponding FQN for each function call. We checked whether the corresponding FQNs matched the APIs in the collected API mappings, identifying the first matched API as the reference API.
We denote the containing function as an outdated function if a deprecated API was matched. Meanwhile, we denote it as an up-to-dated function if a replacing API was matched. In total, we collected Python 113,660 source files by querying SourceGraph. Among them, we extracted 9,022 outdated and 19,103 up-to-dated functions. The statistics of the outdated and up-to-dated functions are presented in Table I.

### v3

# LLMs Meet Library Evolution: Evaluating Deprecated API Usage in LLM-based Code Completion

@@ Abstract
To address this gap, we conducted the first evaluation study on deprecated API usage in LLM-based code completion. This study involved seven advanced LLMs, 145 API mappings from eight popular Python libraries, and 28,125 completion prompts. The study results reveal the status quo (i.e., API usage plausibility and deprecated usage rate) of deprecated API and replacing API usage in LLM-based code completion from the perspectives of model, prompt, and library, and indicate the root causes behind. Based on these findings, we propose two lightweight fixing approaches, ReplaceAPI and InsertPrompt, which can serve as baseline approaches for future research on mitigating deprecated API usage in LLM-based completion. Additionally, we provide implications for future research on integrating library evolution with LLM-driven software development.
@@ I Introduction
Large language models (LLMs) [1, 2, 3, 4, 5, 6] have significantly advanced various aspects of software engineering [7], including code completion [8, 9, 10], code understanding [11, 12], defect detection [13, 14], and program repair [15, 16, 17]. These models, pre-trained or fine-tuned with extensive knowledge of code on large corpora, are effective for tailoring to different downstream tasks. In the realm of code completion, the state-of-the-art has evolved from statistics-based methods [18, 19] to LLM-based techniques [20, 21, 22]. Code completion is a sophisticated task that suggests variables, functions, classes, methods, and even entire code blocks, which depends on developers’ practical needs.
@@
Motivation. To accelerate development, developers heavily rely on third-party libraries, interacting with them through Application Programming Interfaces (APIs). However, this reliance presents a challenge for code completion tools. Third-party libraries constantly evolve to undergo refactorings [23], fix bugs [24], apply security patches [25], or introduce new features. This rapid evolution leads to frequent API changes, with older APIs being deprecated and replaced by newer ones. Deprecated APIs are discouraged to use because of their incompatiblity with newer features or data, which will eventually disappear in future library updates [26]. Taking PyTorch [27], a popular deep learning library for instance, the API torch.gels() was deprecated in version 1.2 (August 2019) in favor of torch.lstsq(). Then, torch.lstsq() was deprecated in version 1.9 (June 2021) in favor of torch.linalg.lstsq(). Consequently, newly developed code should avoid using the deprecated torch.gels() and torch.lstsq(). Therefore, it’s crucial for code completion tools to suggest correct and up-to-dated APIs to developers.
@@
Literature. However, to the best of our knowledge, the capabilities of LLM-based code completion regarding API deprecation is understudied [28, 29]. Although there emerges a substantial number of evaluation on code completion, a body of research focused on assessing the overall accuracy across various benchmarks [11, 30, 31, 32, 33]. Interestingly, Ding et al. [34] identified undefined names and unused variables as the most common syntactic errors produced by LLMs in Python code completions. Izadi et al. [35] found that incorrect function name predictions were prevalent, accounting for 23% of all token-level errors. Furthermore, Recent studies [36, 37] highlighted the issue of hallucinations in LLM-generated code. Their findings indicate the prevalence and potential risks of using unexpected APIs. Nevertheless, while researchers have noted the prevalence of incorrect function name predictions, they have not investigated this issue in depth. Library APIs, which constitute an important part in predicting external function names, are worth attached importance to.
@@
Study. To address this gap, we conducted a study to examine the deprecated API usage in LLM-based code completion. The study aims to answer the primary research question:
@@
What are the status quo (i.e., API Usage Plausibility (AUP) and Deprecated Usage Rate (DUR)) and potential root causes of deprecated and replacing API usage in LLM-based code completion?
@@
Model Perspective (RQ1) investigates the status quo and potential causes based on the performance of various LLMs;
@@
Prompt Perspective (RQ2) examines the impact of different prompts on the status quo and potential causes;
@@
Library Perspective (RQ3) analyzes the status quo and potential causes across different libraries.
@@
To address these research questions, we conducted a series of experiments involving various libraries and LLMs. We collected 145 API mappings between deprecated APIs and their replacements from eight popular Python libraries. Based on these mappings, we retrieved 9,022 outdated functions and 19,103 up-to-dated functions using the deprecated APIs and replacing APIs, respectively. Then, we identify the locating lines of the deprecated or replacing APIs, mask them and subsequent code, and use the remaining parts as the code completion prompts. The original deprecated or replacing API is referred as the reference API. These prompts were then inputted into seven advanced code LLMs, including CodeLlama [5] and GPT-3.5, to generate completions and analyze the predicted API usages. If the predicted API usage corresponds to either the deprecated or replacement version of the reference API, it is annotated as plausible; otherwise, it is annotated as Others.
@@
The study results reveal the following findings: Finding in RQ1: All evaluated LLMs encounter challenges in predicting plausible API usages and face issues with deprecated API usages, due to the presence of deprecated API usages during model training and the absence of API deprecation knowledge during model inference. Finding in RQ2: For the two categories of prompts derived from outdated and up-to-dated functions, the LLMs’ performance in predicting plausible and deprecated API usages differs significantly, influenced by the distinct code context characteristics of these prompts. Finding to RQ3: Across the eight libraries, the LLMs exhibit significant differences in their use of deprecated APIs, influenced by the characteristics of API deprecations during library evolution.
@@
Lightweight Mitigation. Based on the study results and findings, we explored the feasibility of using two lightweight fixing approaches to mitigate deprecated API usage in LLM-based code completion. Given a completion containing a deprecated API usage, the first approach, named ReplaceAPI, directly replaces the deprecated API usage with the replacement and regenerates the remaining parts (e.g., argument list) during the decoding process. The second approach, named InsertPrompt, inserts an additional replacing prompt after the original prompt to guide the LLMs to use the replacing API and then regenerate the completions. We then evaluate the effectiveness of the proposed approaches in terms of fixing deprecated API usages and the accuracy in predicting line-level completions (RQ4). The evaluation results demonstrate that ReplaceAPI effectively addresses deprecated API usages for all evaluated open-source LLMs, achieving fix rates exceeding 85% with acceptable accuracy measured by Edit Similarity and Exact Match compared to ground-truth completions. While InsertPrompt does not currently achieve sufficient effectiveness and accuracy in fixing completions containing deprecated API usage, it shows potential for future exploration.
Contribution. This paper makes the following contributions:
  • We conducted the first study that reveals the status quo and causes of deprecated API usages in LLM-based code completion from prespectives of model, prompt, and library.
@@
  • We proposed two lightweight approaches, named ReplaceAPI and InsertPrompt, to serve as baselines for mitigating deprecated API usage in LLM-based completion.
@@ II-A Library Evolution
Library evolution involves refactorings [23], bug fixes [24], and new feature introductions. Typically, refactorings can deprecate old APIs and introduce new replacements. Several studies have examined the reasons that developers deprecate APIs and how the clients react to such deprecations [38, 39, 40, 41]. The reasons include improving readability, reducing redundancy, avoiding bad code practices and fixing functional bugs. Deprecated APIs can affect hundreds of clients [42], particularly when clients struggle to keep pace with rapidly evolving software [43, 44]. McDonnell et al. [45] found that only 22% of outdated API usages are eventually upgraded to use replacing APIs. Similarly, Hora et al. [46] found that client developers consumes considerate time to discover and apply replacing APIs, with the majority of systems not reacting at all. When clients do not upgrade their APIs, they silently accumulate technical debt in the form of future API changes when they finally upgrade [47]. To locate the replacing API, existing works leverage change rules written by developers [48], developer recordings [49], similarity matching [50], mining API usage in libraries [51], and in client projects [52]. Henkel and Diwan [49] developed an IDE plugin that allows library developers to record API refactoring actions and client developers to replay them. Godfrey and Zou [53] proposed a semi-automated origin analysis using similarities in name, declaration, complexity metrics, and call dependencies. Wu et al. [54] introduced a hybrid approach combining call dependency and text similarity analysis to identify API change rules. Recently, [55] proposed RepFinder to find replacing APIs for deprecated APIs in library updates from multiple sources.
In this work, we aim to comprehend the statuses and causes of deprecated API usage in LLM-based code completion and provide implications for mitigating the deprecated API usages.
@@ II-B LLM-based Code Completion
Code completion is an important functionality in modern IDEs and editors. Historically, researchers have explored statistical models [18, 19]. With the advent in natural language processing, researchers have embraced deep learning for code completion [8, 9] because they are similar in token-based prediction. To explore the capability of code completion tools driven by LLMs [20, 21, 22], numerous evaluations of LLMs have been proposed. Ciniselli et al. [32, 33] conducted a large-scale study exploring the accuracies of state-of-the-art Transformer-based models in supporting code completion at various granularity levels, from single tokens to entire code segments. Zeng et al. [11] found that pre-trained models significantly outperform non-pre-trained state-of-the-art techniques in program understanding tasks. They also reveal that no single pre-trained model dominates across all tasks. Xu et al. [30] evaluated the performance of LLMs on the HumanEval dataset. Ding et al. [34] identified undefined names and unused variables as the most common errors produced by language models in Python code completions. Izadi et al. [35] evaluated the LLMs using real auto-completion usage data across 12 languages. They found that incorrect function name predictions, were prevalent, accounting for 23% of all token-level errors. Besides, Liu et al. [31] proposed EvaluPlus, which benchmarks the functional correctness of LLM-synthesized code using test cases. In addition to accuracy concerns, LLM-based approaches face issues such as security vulnerabilities and hallucinations. Sallou et al. [56] explored threats posed by LLMs, including unpredictability in model evolution, data leakage, and reproducibility. Liu et al. [36] categorized the hallucinations brought by LLM-generated code.
The findings on incorrect function predictions partially motivate our study. However, we focus on the severity of predicting deprecated API usages in LLM-based code completion.
@@ III Study Setup
We chose Python, a popular programming language which ranks first among the most popular ones based in the recent year [57]. We targeted eight popular Python libraries. Five of these libraries were used in a previous study on Python API deprecation [25], including Numpy, Pandas, scikit-learn, SciPy, and seaborn. Additionally, we added three popular deep learning libraries, i.e., TensorFlow, PyTorch, and Transformers. The setup of our study is presented in Figure 1. It includes four steps. The API Mapping Collection gathers mappings between deprecated APIs and their replacements from various libraries. The Completion Prompt Construction step involves creating completion prompts by identifying instances of deprecated and replacing API usage in open-source Python repositories. The LLM-Based Code Completion step uses various LLMs to generate code completions for these prompts. Finally, the Completion Result Annotation step automatically annotates the generated completions and calculates relevant metrics.
@@ III-A API Mapping Collection
We identified API mappings (i.e., deprecated APIs and the mapping replacements) from the documentation and change logs from each library version following the previous study [25]. For each library, we selected their versions that were released after January 2019. We chose January 2019 as the starting point for collecting API mappings, providing a five-year window that balances sufficient version numbers with manageable human effort. The version ranges are presented in Table I. Specifically, we reviewed the documentation and change logs of each library and manually look for deprecated API occurrences which indicate the corresponding the mapping replacements. For instance, in the API documentation of PyTorch, version 1.9.0 [58], a deprecation message indicates that “torch.lstsq() is deprecated in favor of torch.linalg.lstsq() and will be removed in a future PyTorch release.”, where the mapping of the deprecated API to the replacing API is torch.lstsq $\to$ torch.linalg.lstsq. For one-to-many mappings (i.e., one deprecated API mapped to multiple replacing APIs), we split them into many one-to-one mappings. Two authors independently collected the data, with a third author resolving inconsistencies. The process took about three working days, and the Jaccard coefficient between the first two collectors’ mappings was 91.8%. In total, we collected a preliminary number of 247 API mappings.
TABLE I: Statistics of our Collected API Mappings
  Library Version Range # Mappings # Functions
@@
  Numpy 1.16.0-1.26.4 3 567 2,988
Pandas 0.24.0-2.2.2 10 69 69
scikit-learn 0.21.3-1.5.0 18 985 1,197
SciPy 1.2.1-1.13.0 4 245 1,458
seaborn 0.9.1-0.13.2 3 904 1,329
TensorFlow 1.13.1-2.16.1 57 1,491 4,830
PyTorch 1.0.1-2.3.0 21 4,726 6,406
Transformers 1.0.0-4.40.2 29 100 63
@@ III-B Completion Prompt Construction
We constructed code completion prompts by searching real-world code snippets that contain usages of either deprecated APIs or replacing APIs from collected API mappings in open-source Python repositories, to simulate realistic code completion scenarios.
@@
We utilized Sourcegraph [59], a widely used code search service to search code snippets. It supports integration with GitHub where we can retrieve Python source files from millions of open-source code repositories. For each deprecated or replacing API, we constructed search queries using both its full qualified name (FQN) (e.g., torch.lstsq) and a logical disjunction of its constituent parts (e.g., “torch AND linalg AND lstsq”) to ensure comprehensive retrieval. For each retrieved Python source file, we parsed it into an Abstract Syntax Tree (AST) and extracted the containing functions that invoked the deprecated or replacing APIs. Specifically, we located function definition nodes in the AST and traversed its descendants. For each descendant, we checked if it is a function call node and matched the function call to the deprecated or replacing APIs. To correctly match the function call via API FQNs, we performed lightweight object type resolution and alias resolution, similar to [25].
@@
  • Alias Resolution: Developers can alias packages, classes, and functions in Python using the import-as feature [60]. This mechanism requires resolving API aliases by analyzing import statements. For example, the pandas package is often imported with the alias “pd” via the statement import pandas as pd. In this case, pd.DataFrame.loc() was resolved to pandas.DataFrame.loc(). Additionally, Python provides a from-import mechanism allowing developers to use APIs with short names instead of their FQNs. For example, through from torch.linalg import lstsq, the API in torch.linalg can be directly called via lstsq(). These short names were resolved by analyzing the import statements.
After the lightweight object type resolution and alias resolution, we obtained the corresponding FQN for each function call. We checked whether the corresponding FQNs matched the APIs in the collected API mappings, identifying the first matched API as the reference API.
We denote the containing function as an outdated function if a deprecated API was matched. Meanwhile, we denote it as an up-to-dated function if a replacing API was matched. We collected 113,660 Python source files by querying SourceGraph. We filtered out API mappings that returned either no instances of outdated or up-to-dated functions. As a result, we collected 9,022 outdated and 19,103 up-to-dated functions from 145 API mappings. The statistics are presented in Table I.